# PaCE: Parsimonious Concept Engineering for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are being used for a wide variety of tasks. While they are capable of generating human-like responses, they can also produce undesirable output including potentially harmful information, racist or sexist language, and hallucinations. Alignment methods are designed to reduce such undesirable output, via techniques such as fine-tuning, prompt engineering, and representation engineering. However, existing methods face several challenges: some require costly fine-tuning for every alignment task; some do not adequately remove undesirable concepts, failing alignment; some remove benign concepts, lowering the linguistic capabilities of LLMs. To address these issues, we propose Parsimonious Concept Engineering (PaCE), a novel activation engineering framework for alignment. First, to sufficiently model the concepts, we construct a large-scale concept dictionary in the activation space, in which each atom corresponds to a semantic concept. Given any alignment task, we instruct a concept partitioner to efficiently annotate the concepts as benign or undesirable. Then, at inference time, we decompose the LLM activations along the concept dictionary via sparse coding, to accurately represent the activation as a linear combination of the benign and undesirable components. By removing the latter ones from the activation, we reorient the behavior of LLMs towards alignment goals. We conduct experiments on tasks such as response detoxification, faithfulness enhancement, and sentiment revising, and show that PaCE achieves state-of-the-art alignment performance while maintaining linguistic capabilities. Our collected dataset for concept representations is available at https://github.com/peterljq/Parsimonious-Concept-Engineering.
@@ 1 Introduction
Large Language Models (LLMs) are useful for tasks as far ranging as question answering [76, 65], symbolic reasoning [57, 25], multi-modal synthesis [86, 41, 45], and medical diagnosis [85]. LLMs are typically pre-trained on a broad collection of textual corpora with the next-token prediction objective [55, 70], enabling them to generate human-like text. An important aspect of deploying pre-trained LLMs for real-world applications is preventing undesirable responses such as toxic language, hallucinations, and biased information through alignment methods, which aim to make AI systems behave in line with human intentions and values [28]. A common alignment approach is tuning LLMs with human feedback [56, 62] for better instruction-following capabilities. However, after such aligning, undesirable and harmful content can still be elicited from LLMs. For example, jailbreaking can produce hate speech and aggression [22, 32], stress-testing shows hallucinatory responses such as illogical statements [87], and various kinds of bias are not fully removed from LLM responses [19]. This emphasizes the need for further development towards aligned LLMs.
@@
  • Concept Dictionary Construction and Partition (§3.2): Since existing works only provide a limited number of concept directions, we collect a large concept dictionary, PaCE-1M, that consists of 40,000 concept directions extracted from over 1,200,000 context sentences. In particular, for each concept in the Brown Corpus [18], we use a knowledge-driven GPT [65, 45, 36] to propose contextual scenarios to describe the concept, and extract concept directions in the representation (activation) space [88] from the context sentences. This is done only once offline. Given an alignment task, we further instruct a GPT to automatically partition the concept directions in the dictionary into benign and undesirable directions.
@@
We evaluate PaCE on multiple alignment tasks including response detoxification, faithfulness enhancement, and sentiment revising (§4). We show that PaCE achieves state-of-the-art performance on these tasks, while retaining its linguistic capability at a comparable level. We further shed insights on the concept directions of PaCE-1M by showing that they are geometrically consistent with their concept semantics and a decomposition reveals the semantics of the activations.
@@ 2.1 The Latent Space and Its Linear Controllability
A classic setting where linear controllability shows up is that of word embeddings. Here, ${\mathcal{T}}$ is the vocabulary (say, the set of English words), ${\mathcal{Z}}$ contains some vectors in ${\mathbb{R}}^{d}$, and $g$ is a bijection between ${\mathcal{Z}}$ and ${\mathcal{T}}$. In the seminal work of Mikolov et al. [51], the authors observe that word embeddings learned by recurrent neural networks approximately enjoy relations such as ${\boldsymbol{z}}_{\text{king}}-{\boldsymbol{z}}_{\text{man}}+{\boldsymbol{z}}_{\text{woman}}\approx{\boldsymbol{z}}_{\text{queen}}$, where one can view ${\boldsymbol{z}}_{\text{woman}}-{\boldsymbol{z}}_{\text{man}}$ as the concept direction ${\boldsymbol{v}}\in{\mathcal{V}}$ and the control strength to be $c=1$. This observation is later extended to word embeddings of various networks and learning objectives such as word2vec [50], Skip-Grams [49, 35], GloVe [59], and Swivel [63]. On the theoretical front, a fruitful line of research has been devoted to understanding the emergence of such properties in word embeddings [3, 21, 2, 1, 17, 53].
@@
Modern neural architectures such as transformers have significantly boosted the linguistic performance of language models. Much of their success is attributed to the attention mechanism, which incorporates long-range context into the neural activations in transformers. This has motivated people to take ${\mathcal{Z}}$ as certain hidden states in transformers [^2] A variety of choices of layers have been explored in the literature; see, e.g., [67] for a comparison., and search for concept directions ${\mathcal{V}}$. This has led to a fascinating line of works supporting the empirical existence of ${\mathcal{V}}$: [6, 47] find directions that indicate truthful output, [68] finds directions for sentiments, [88] finds directions for emotions and honesty, and [54] finds directions for current player tile in a synthetic board game model. Interestingly, [58, 30] further offer theoretical models, under which the linear controllability shows up provably in the latent space of LLMs.
@@ 2.2 Controlling Language Models via Latent Space Engineering
where $\hat{c}>0$ is a parameter of the strength of control. In principle, as the input prompt may contain a different ‘extent’ of the concept to be removed, $\hat{c}$ should depend on both the input prompt and the concept. Thus, in practice, one either tunes $\hat{c}$ per input prompt and concept, which is laborious, or one fixes a $\hat{c}$, which is sub-optimal. Indeed, this has been observed by the work [71]: In their Table 10, the optimal coefficients $\hat{c}$ are markedly different across the examples; see also their ‘discussion’ section.
@@
Here, for any ${\boldsymbol{w}}\in{\mathbb{R}}^{d}$, $\operatorname{span}({\boldsymbol{w}})$ is the linear subspace spanned by ${\boldsymbol{w}}$, and for any linear subspace ${\mathcal{S}}\subset{\mathbb{R}}^{d}$, $\Pi_{{\mathcal{S}}}$ denotes the ortho-projector onto ${\mathcal{S}}$. Such an idea is later applied to neural activations of LLMs [23, 88]. Applying orthogonal projection to remove concept directions from latent codes may be reasonable: if directions corresponding to different concepts are orthogonal, then orthogonal projection does not remove directions from concepts other than the gender direction. That being said, there are often more concept directions presented, and they are not orthogonal. For example, [29] shows that causally related concepts only exhibit partial orthogonality for their directions.
@@ 3.3 Overcomplete Oblique Projection via Sparse Coding
    where $\Omega({\boldsymbol{c}})$ is a sparsity-promoting regularizer that we will discuss soon. Then, each coefficient $c_{i}^{\text{in}}$ for $i\in\{1,\dots,n\}$ can be viewed as how much the concept represented by ${\boldsymbol{v}}_{i}$ is in ${\boldsymbol{z}}^{\text{in}}$, and ${\boldsymbol{r}}^{\text{in}}$ is the residual that is not explained by ${\boldsymbol{D}}$.
@@
The synthesized ${\boldsymbol{z}}^{\text{ctrl}}$ will replace ${\boldsymbol{z}}^{\text{in}}$ to be passed on to the next layer of the neural network.
@@ 4.1 Improving Safety by Response Detoxification
Safety Responses. The evaluation has nine categories: Political Sensitivity (PS), Pornography (PG), Ethics and Morality (EM), Illegal Activities (IA), Mental Harm (MH), Offensiveness (OF), Physical Harm (PH), Privacy and Property (PP), and Unfairness & Bias (UB). As shown in Table 1, for LLaMa2-7B, PaCE improves by 60-80% over the vanilla method in categories including IA, MH, OF, PH, PP, and UB. When compared to other methods, PaCE performs competitively and improves by 6-20%. While our method did not perform the best in PS, PG, and EM, the gap for those categories is relatively small considering the significant overall gains. Notably, for LLaMA2-13B which has more parameters and a presumably more structured latent space, PaCE dominates other methods in all categories, demonstrating the necessity for respecting the latent structures when modifying representations. Finally, Table 3 shows the contribution of design choices in PaCE, and Table 3 shows the effect of the dictionary size on the performance. We observe clear improvement after each design choice is progressively added to PaCE. Appendix B.5 includes the details of these ablation studies.
@@
4.2 Improving Faithfulness and Removing Negative
Sentiment
@@ 4.2 Improving Faithfulness and Removing Negative
Setup. Faithfulness: We use the FactScore suite and the fact evaluator for faithful biography generation [52]. The suite is divided into labeled and unlabeled subsets used in different sections of the original paper. Our table reports the Labeled Score (LS), the total number of Labeled Atomic Facts (LAF), Unlabeled Score (US), and the total number of unlabeled Atomic Facts (LAF). Sentiment: We use the HolisticBias suite [66] and hate speech evaluator [64] to measure the sentiment of the response to underrepresented descriptors. The reported numbers are the average of non-negative sentiment scores for underrepresented groups categorized by Gender (GN), Occupation (OC), and Nationality (NT). During the sentiment revising, the concept setups for all approaches follow the detoxification setup. For the faithfulness experiments, PaCE removes the top 50 undesirable (hallucinatory) concepts ranked by the GPT partitioner. The Prompting approach instructs the LLM not to output sentences relevant to these top concepts. The VecAdd and OrthoProj approaches operate on the concept vector of ‘fabrication’.
Figure 7: Concept coefficients solved by PaCE are an interpretable interface, and they are further used for activation intervention.
@@ 4.3 Representation Space Sampled by PaCE-1M
Interpretability. Concept bottlenecks, where the model input is mapped to a list of human-readable concepts for interpretable decisions, are widely adopted for intervening model behaviors [34, 80, 7]. PaCE’s decomposition of the input prompt onto the large-scale concept dictionary also enables us to investigate the LLM’s internal behavior regarding the input prompt. Figure 7 shows the PaCE-solved weights for top nice concepts (in the order of absolute magnitude) in the activation space for an input prompt. The decomposition indicates that the target LLM corresponds the prompt toward the concept ‘irrational’ and against ‘conclusive’, which enables PaCE to execute the followed-up activation intervention (e.g., remove the concept ‘irrational’ by §3.3).
@@
5 Conclusion and Discussion
@@
This research was supported by ARO MURI contract W911NF-17-1-0304, DARPA GARD HR001119S0026, DARPA RED HR00112090132, ODNI IARPA HIATUS Program Contract #2022-22072200005, and NSF Graduate Research Fellowship #DGE2139757. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.

### v2

@@ Abstract
Large Language Models (LLMs) are being used for a wide variety of tasks. While they are capable of generating human-like responses, they can also produce undesirable output including potentially harmful information, racist or sexist language, and hallucinations. Alignment methods are designed to reduce such undesirable outputs via techniques such as fine-tuning, prompt engineering, and representation engineering. However, existing methods face several challenges: some require costly fine-tuning for every alignment task; some do not adequately remove undesirable concepts, failing alignment; some remove benign concepts, lowering the linguistic capabilities of LLMs. To address these issues, we propose Parsimonious Concept Engineering (PaCE), a novel activation engineering framework for alignment. First, to sufficiently model the concepts, we construct a large-scale concept dictionary in the activation space, in which each atom corresponds to a semantic concept. Given any alignment task, we instruct a concept partitioner to efficiently annotate the concepts as benign or undesirable. Then, at inference time, we decompose the LLM activations along the concept dictionary via sparse coding, to accurately represent the activations as linear combinations of benign and undesirable components. By removing the latter ones from the activations, we reorient the behavior of the LLM towards the alignment goal. We conduct experiments on tasks such as response detoxification, faithfulness enhancement, and sentiment revising, and show that PaCE achieves state-of-the-art alignment performance while maintaining linguistic capabilities. Our collected dataset for concept representations is available at https://github.com/peterljq/Parsimonious-Concept-Engineering.
@@ 1 Introduction
Large Language Models (LLMs) are useful for tasks as far ranging as question answering [77, 65], symbolic reasoning [56, 25], multi-modal synthesis [86, 40, 44], and medical diagnosis [85]. LLMs are typically pre-trained on a broad collection of textual corpora with the next-token prediction objective [54, 70], enabling them to generate human-like text. An important aspect of deploying pre-trained LLMs for real-world applications is preventing undesirable responses such as toxic language, hallucinations, and biased information through alignment methods, which aim to make AI systems behave in line with human intentions and values [28]. A common alignment approach is tuning LLMs with human feedback [55, 62] for better instruction-following capabilities. However, after such aligning, undesirable and harmful content can still be elicited from LLMs. For example, jailbreaking can produce hate speech and aggression [22, 32], stress-testing shows hallucinatory responses such as illogical statements [87], and various kinds of biases are not fully removed from LLM responses [19]. This emphasizes the need for further development of aligned LLMs.
@@
  • Concept Dictionary Construction and Partition (§3.2): Since existing works only provide a limited number of concept directions, we collect a large concept dictionary, PaCE-1M, that consists of 40,000 concept directions extracted from over 1,200,000 context sentences. In particular, for each concept in the Brown Corpus [18], we use a knowledge-driven GPT [65, 44, 35] to propose contextual scenarios to describe the concept, and extract concept directions in the representation (activation) space [88] from the context sentences. This is done only once offline. Further, given any alignment task, we instruct a GPT to automatically partition the concept directions in the dictionary into benign and undesirable directions, which is done once per task offline.
@@
We evaluate PaCE on alignment tasks including response detoxification, faithfulness enhancement, and sentiment revising (§4). We show that PaCE achieves state-of-the-art performance on these tasks, while retaining its linguistic capability at a comparable level. We further shed insights on the concept directions of PaCE-1M: concept directions tend to form clusters with directions from each cluster corresponding to similar semantics, and decomposing an activation reveals its semantics.
@@ 2.1 The Latent Space and Its Linear Controllability
A classic setting where linear controllability shows up is that of word embeddings. Here, ${\mathcal{T}}$ is the vocabulary (say, the set of English words), ${\mathcal{Z}}$ contains some vectors in ${\mathbb{R}}^{d}$, and $g$ is a bijection between ${\mathcal{Z}}$ and ${\mathcal{T}}$. In the seminal work of Mikolov et al. [50], the authors observe that word embeddings learned by recurrent neural networks approximately enjoy relations such as ${\boldsymbol{z}}_{\text{king}}-{\boldsymbol{z}}_{\text{man}}+{\boldsymbol{z}}_{\text{woman}}\approx{\boldsymbol{z}}_{\text{queen}}$, where one can view ${\boldsymbol{z}}_{\text{woman}}-{\boldsymbol{z}}_{\text{man}}$ as the concept direction ${\boldsymbol{v}}\in{\mathcal{V}}$ and $c=1$ as the control strength. This observation is later extended to word embeddings of various networks and learning objectives such as word2vec [49], Skip-Grams [48, 34], GloVe [59], and Swivel [63]. On the theoretical front, a fruitful line of research has been devoted to understanding the emergence of such properties in word embeddings [3, 21, 2, 1, 17, 52].
@@
Modern neural architectures such as transformers have significantly boosted the linguistic performance of language models. Much of their success is attributed to the attention mechanism, which incorporates long-range context into the neural activations in transformers. This has motivated people to take ${\mathcal{Z}}$ as certain hidden states in transformers [^2] A variety of choices of layers have been explored in the literature; see, e.g., [67] for a comparison., and search for concept directions ${\mathcal{V}}$ in ${\mathcal{Z}}$. An interesting line of works has supported the empirical existence of ${\mathcal{V}}$: [6, 46] find directions that indicate truthful output, [68] finds directions for sentiments, [88] finds directions for emotions and honesty, and [53] finds directions for current player tile in a synthetic board game model. Interestingly, [71, 57, 30] further offer theoretical models, under which the linear controllability shows up provably in the latent space of LLMs.
@@ 2.2 Controlling Language Models via Latent Space Engineering
where $\hat{c}>0$ is a parameter of the control strength. In principle, as each input prompt may contain a different ‘extent’ of the concept to be removed, $\hat{c}$ should depend on both the prompt and the concept. Thus, in practice, one either tunes $\hat{c}$ per input prompt and concept, which is laborious, or one fixes a $\hat{c}$, which is sub-optimal. Indeed, this has been observed by the work [72]: In their Table 10, the optimal coefficients $\hat{c}$ are markedly different across the examples; see also their ‘discussion’ section.
@@
Here, for any ${\boldsymbol{w}}\in{\mathbb{R}}^{d}$, $\operatorname{span}({\boldsymbol{w}})$ is the linear subspace spanned by ${\boldsymbol{w}}$, and for any linear subspace ${\mathcal{S}}\subset{\mathbb{R}}^{d}$, $\Pi_{{\mathcal{S}}}$ denotes the ortho-projector onto ${\mathcal{S}}$. Such an idea is later applied to neural activations of LLMs [23, 88]. Applying orthogonal projection to remove concept directions from latent codes may be reasonable: if directions corresponding to different concepts are orthogonal, then orthogonal projection only removes the gender direction while leaving the others intact. That being said, there are often more concept directions presented, and they are not orthogonal. For example, [29] shows that causally related concepts only exhibit partial orthogonality for their directions.
@@ 3.3 Overcomplete Oblique Projection via Sparse Coding
    where $\Omega({\boldsymbol{c}})$ is a sparsity-promoting regularizer that we will discuss soon. Then, each coefficient $c_{i}^{\text{in}}$ for $i\in\{1,\dots,n\}$ can be viewed as how much the concept represented by ${\boldsymbol{v}}_{i}$ is present in ${\boldsymbol{z}}^{\text{in}}$, and ${\boldsymbol{r}}^{\text{in}}$ is the residual that is not explained by ${\boldsymbol{D}}$.
@@
The synthesized ${\boldsymbol{z}}^{\text{ctrl}}$ will replace ${\boldsymbol{z}}^{\text{in}}$ to be passed on to the next layer of the neural network [^4] While in sparse coding one typically removes the residual ${\boldsymbol{r}}^{\text{in}}$ for denoising purpose, here ${\boldsymbol{r}}^{\text{in}}$ may contain useful information (e.g., grammar) not captured by the dictionary, so we simply keep it in the synthesized ${\boldsymbol{z}}^{\text{ctrl}}$..
@@ 4.1 Improving Safety by Response Detoxification
Safety Responses. The evaluation has nine categories: Political Sensitivity (PS), Pornography (PG), Ethics and Morality (EM), Illegal Activities (IA), Mental Harm (MH), Offensiveness (OF), Physical Harm (PH), Privacy and Property (PP), and Unfairness & Bias (UB). Table 1 shows that, for LLaMa2-7B, PaCE improves by 60-80% over the vanilla method in categories including IA, MH, OF, PH, PP, and UB. When compared to other methods, PaCE performs competitively and improves by 6-20%. While our method did not perform the best in PS, PG, and EM, the gap for those categories is relatively small considering the significant overall gains. Notably, for LLaMA2-13B which has more parameters and a presumably more structured latent space, PaCE dominates other methods in all categories, demonstrating the necessity for respecting the latent structures when modifying representations. Finally, Table 3 shows the contribution of design choices in PaCE, and Figure 6 shows the effect of the dictionary size on the performance. We observe clear improvement after each design choice is progressively added. Appendix B.5 includes the details of these ablation studies.
@@
Solvers. Figure 7 additionally evaluates Orthogonal Matching Pursuit (OMP) [58, 7], a fast greedy solver for the activation decomposition. OMP iteratively adds to the support the concept that has maximum coherence with the unexplained residual and updates the residual by solving the least square using the new support. It stops when a pre-defined maximum size $k$ of support is reached. Intuitively, the $k$ is the number of non-zero elements in the solved coefficients. We observe from the table that one can choose improvements in computational speed at the cost of safety performance.
@@
4.2 Improving Faithfulness and Removing Negative Sentiment
@@
Figure 7: Ablation study for solvers. We observe that greedy solvers can improve computational speed at the cost of safety performance.
Method Time per decomposition (s, $\downarrow$) Safety (%, $\uparrow$)
OMP ($k=50$) 0.045 63.1
OMP ($k=100$) 0.182 64.4
OMP ($k=150$) 0.381 66.9
OMP ($k=200$) 0.749 70.8
Elastic Net 0.411 72.0
Setup. Faithfulness: We use the FactScore suite and the fact evaluator for faithful biography generation [51]. The suite is divided into labeled and unlabeled subsets used in different sections of the original paper. Our table reports the Labeled Score (LS), the total number of Labeled Atomic Facts (LAF), the Unlabeled Score (US), and the total number of unlabeled Atomic Facts (LAF). Sentiment: We use the HolisticBias suite [66] and hate speech evaluator [64] to measure the sentiment of the response to underrepresented descriptors. The reported numbers are the average of non-negative sentiment scores for underrepresented groups categorized by Gender (GN), Occupation (OC), and Nationality (NT). During the sentiment revising, the concept setups for all approaches follow the detoxification setup. For the faithfulness experiments, PaCE removes the top 50 undesirable (hallucinatory) concepts ranked by the partitioner. The Prompting approach instructs the LLM not to output sentences relevant to these top concepts. The VecAdd and OrthoProj operate on the concept vector of ‘fabrication’.
@@ 4.3 Representation Space Sampled by PaCE-1M
5 Discussion
We provide discussions on the monosemanticity of concepts and connections among different alignment paradigms in this section. We also argue how PaCE handles context-dependent concepts.
5.1 Polysemy of Words
While VecAdd and OrthoProj may be affected by the polysemy of words, PaCE’s overcomplete dictionary allows accurate analysis of the target representation through sparse decomposition. Table 1 and Table 4 show that PaCE outperforms OrthoProj and VecAdd on linguistic metrics. We attribute the high helpfulness of PaCE to the large-scale dictionary with sparse coding, explained as follows.
Comprehensive Coverage. Since the dictionary is large, concepts with single and clear semantics are involved. E.g., if the stimuli of ‘kill’ may have different meanings, there exist other more polarized concept vectors such as ‘murder’ (more harmful) and ‘spend’ (more benign).
Parsimony of Solution. Sparse coding aims to choose the fewest concepts to reconstruct the latent representation (i.e., parsimony). For the sake of argument, assuming the sentence is about ‘killing time’ and the vanilla LLM has the correct semantic understanding of its benignness, the latent representation of the whole sentence will be closer to concepts such as ‘spend’ and ‘time’ rather than string-matching to ‘kill’ (which in your setup could have mixed harmful and benign senses). As the sparse coding of the target representation promotes the parsimonious selection of concepts with monosemantics, it helps to represent benign contexts correctly without assigning significant weights to ambiguous terms like ‘kill’.
5.2 Different Alignment Paradigms
As mentioned in §1, beyond activation engineering, there are other alignment paradigms such as Supervised Fine-Tuning (SFT) [26], Reinforcement Learning from Human Feedback (RLHF) [55], and Knowledge Engineering (KE) [14, 74]. We clarify the main advantages of PaCE over them.
Training-Free. RLHF, SFT, and KE all need to tune the parameters of LLM, which potentially degrade the well-structured priors of the pre-trained LLM. Taking a step back, even if LoRA is adopted for these paradigms, the training/tuning incurs significant computation and memory costs. PaCE does not modify the parameters of LLM and requires no training. It better preserves the priors of LLM, provides a low-resource alignment solution, and retains the general linguistic capabilities.
Interpretable and Adaptive. The solved coefficients of PaCE are an accurate interpretation of how a user input’s representation is composed in the concept space. Also, when a new alignment goal is set, RLHF, SFT, and KE need to collect sufficient task samples and tune the LLM on the new dataset. In contrast, PaCE just needs to run the concept partitioner through CORD, which is expected to be much faster and more convenient.
5.3 Context-dependent Concepts
The structured activation space of LLMs and the large-scale concept dictionary of PaCE help to handle the influence of the context for a concept in the target prompt. As the LLM scales up, its capability to capture and utilize contextual information grows with the help of attention modules. The activation space, as already used for many representation manipulation methods, is expected to convey the underlying semantic information of concepts in the sentence (context). That is, the space hosting concept vectors is not collapsed, and it is structured to distinguish different concepts. The representation (activation) to be steered at inference time encodes the context and conveys the semantics of a concept based on the context. Then, since our overcomplete concept dictionary in PaCE widely covers concepts of various categories, the sparse coding on this dictionary will effectively analyze the target representation as the linear combination of these concepts.
6 Conclusion
@@
This research was supported by ARO MURI W911NF-17-1-0304, DARPA GARD HR001119S0026, DARPA RED HR00112090132, ODNI IARPA HIATUS #2022-22072200005, the NSF grant 2031985, Simons Foundation MoDL 135615, a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI, and NSF Graduate Research Fellowship #DGE2139757. We would like to thank Liangzu Peng, Hancheng Min, Bowen Li, Xinyu Yang, and Fengrui Tian for their suggestions in the presentation and experiments. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, NSF, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.